OpenStudy (idealist10):

Find an integrating factor of the form u(x, y)=P(x)Q(y) for the following equation and solve it.

OpenStudy (idealist10):

\[(\alpha*y+\gamma*x*y)dx+(\beta*x+\delta*x*y)dy=0\]

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Are \(\alpha,\beta,\gamma,\delta\) constants?

OpenStudy (anonymous):

If they are constants, working through the method I used last time gets an integrating factor of \(u(x,y)=\dfrac{1}{xy}\).

OpenStudy (idealist10):

I did use your method but I got stucked in the middle of finding the integrating factor.

OpenStudy (idealist10):

\[F(x)(\beta*x+\delta*x*y)-G(y)(\alpha*y+\gamma*x*y)\]

OpenStudy (anonymous):

Okay, if you recall we have this equation: \[(1)\quad\quad F(x)\underbrace{(\beta x+\delta xy)}_{N(x,y)}-G(y)\underbrace{(\alpha y+\gamma xy)}_{M(x,y)}=M_y-N_x\] where the differential equation was \[M(x,y)~dx+N(x,y)~dy=0\] \(F(x)\) and \(G(y)\) are equivalent to \(\dfrac{u_x}{u}\) and \(\dfrac{u_y}{u}\), respectively, and \(M_y\) and \(N_x\) denote the partial derivatives with respect to the subscript variables. Equation (1) becomes \[\begin{align*}F(x)(\beta x+\delta xy)-G(y)(\alpha y+\gamma xy)&=(\alpha+\gamma x)-(\beta+\delta y)\\\\ F(x)(\beta x+\delta xy)&=G(y)(\alpha y+\gamma xy)+(\alpha+\gamma x)-(\beta+\delta y)\\\\ F(x)&=\frac{G(y)(\alpha y+\gamma xy)+(\alpha+\gamma x)-(\beta+\delta y)}{x(\beta +\delta y)}\end{align*}\] (Note: Solving for \(G(y)\) should get the same result.) As before, since \(F(x)\) is supposed to be a function of \(x\) only, we want to be able to remove any factors of \(y\). To do this, we'll say the numerator is equal to a constant times the function of \(y\) in the denominator. \[G(y)(\alpha y+\gamma xy)+(\alpha+\gamma x)-(\beta+\delta y)=C(\beta+\delta y)\] If we let \(C=-1\), we have a few terms disappear and we can solve for \(G(y)\): \[\begin{align*}G(y)(\alpha y+\gamma xy)+(\alpha+\gamma x)\color{red}{-(\beta+\delta y)}&=\color{red}{-\beta-\delta y}\\\\ G(y)y(\alpha +\gamma x)+(\alpha+\gamma x)&=0\\\\ (G(y)y+1)(\alpha +\gamma x)&=0\\\\ G(y)y+1&=0\\\\ G(y)&=-\frac{1}{y}\end{align*}\] Substituting this into the \(F(x)\) equation gives \[\begin{align*}F(x)&=\frac{-\dfrac{1}{y}(\alpha y+\gamma xy)+(\alpha+\gamma x)-(\beta+\delta y)}{x(\beta +\delta y)}\\\\ F(x)&=\frac{-\alpha -\gamma x+(\alpha+\gamma x)-(\beta+\delta y)}{x(\beta +\delta y)}\\\\ F(x)&=-\frac{\beta+\delta y}{x(\beta +\delta y)}\\\\ F(x)&=-\frac{1}{x}\end{align*}\] In terms of the notation you're given, \(F(x)=P(x)\) and \(G(y)=Q(y)\). The integrating factor is then \[\log u(x,y)=-\int\frac{dx}{x}-\int\frac{dy}{y}~~\implies~~u(x,y)=\frac{1}{xy}\]

OpenStudy (idealist10):

The answer I got is \[\frac{ (\alpha*y+\gamma*x*y)x+(\beta*x+\delta*x*y)y }{ xy }=C\]

OpenStudy (idealist10):

But the book's answer is \[\left| x \right|^{\alpha}*\left| y \right|^{\beta}*e ^{\gamma*x}*e ^{\delta*y}=C\]

OpenStudy (idealist10):

Which is the right answer?

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Distributing the IF gives \[\left(\frac{\alpha}{x}+\gamma\right)dx+\left(\frac{\beta}{y}+\delta\right)dy=0\] Partials: \[\frac{\partial}{\partial y}\left(\frac{\alpha}{x}+\gamma\right)=0\\ \frac{\partial}{\partial x}\left(\frac{\beta}{y}+\delta\right)=0\] Solving for \(\Psi\): \[\begin{align*}\Psi_x&=\frac{\alpha}{x}+\gamma\\\\ \Psi &=\int\left(\frac{\alpha}{x}+\gamma\right)~dx\\\\ &=\alpha\ln|x|+\gamma x+f(y)\\\\ \Psi_y&=f'(y)\\\\ \frac{\beta}{y}+\delta&=f'(y)\\\\ f(y)&=\beta \ln|y|+\delta y+C \end{align*}\] which makes the solution \[\Psi=\alpha\ln|x|+\gamma x+\beta \ln|y|+\delta y=C\] Taking the exponential of both sides will give you the solution from your textbook.

OpenStudy (idealist10):

@SithsAndGiggles , my computer won't show what you typed since processing math is 0%. Can you please take a snapshot of what you just typed?

OpenStudy (idealist10):

Thank you so much, @SithsAndGiggles !!! I'm so sorry that my internet is super slow sometimes.

OpenStudy (anonymous):

You're welcome! And don't worry about it, technology tends to be a inconvenient sometimes.